a shipment of 15 television sets contains four defective ones. in how many ways can one select 5 of these for inspection so that:
a. none of the defective television sets are included
b. both of the defective television sets are included
c. only one of the defective television sets are included
d. either two of the defective television sets or three of the non defective sets are included

Answer by stanbon(75887)

a shipment of 15 television sets contains four defective ones. in how many ways can one select 5 of these for inspection so that:
a. none of the defective television sets are included
11C5 = [11*10*9*8*7]/[1*2*3*4*5] = 462
----------------------------------------------
b. both of the defective television sets are included
Comment: There are four defective; you are selecting 5 TV's; what
is meant by "both"?
-------------------------------------------------
c. only one of the defective television sets are included
Pick a defective: 4 ways
Pick 4 others: 11C4 = 330
Total # of ways: 4*330= 1320
-----------------------------------------------------
d. either two of the defective television sets or three of the non defective sets are included
Those conditions are really the same condition when you are selecting 5.
----------
4C2*11C3 = 990
=========================
Cheers,
Stan H.